40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4431     Entry time: Wed Mar 23 10:34:17 2011
Author: josephb 
Type: Update 
Category: CDS 
Subject: Trend issue fixed 

[Joe, Alex]

Yesterday during the day, Alex ran a script to fix the time stamps in the trends files we had messed up back during the daqd change overs around Feb 17th and 23rd.  See this elog for more information on the trend problem.

Due to how the script runs, basically taking all the data and making a new copy with the correct time stamps, the data collected while the script was running didn't get converted over.  So when he did the final copy of the corrected data, it created a several hour gap in the data from yesterday during the day time.

The original files still exist on the fb machine in /frames/trend/minute_raw_22mar2011 directory.

 

ELOG V3.1.3-